### v1

@@ Appendix B Regression Analysis Additional
Figure B.1: National Comprehensive Cancer Network guidelines example from 2020.
@@
Table B.1 shows the distribution of patient characteristics by optimal care status defined as NCCN treatment adherence.
Table B.1: Sociodemographic and Clinical Characteristics of Massachusetts Cancer Registry Endometrial Cancer Cases from 2015-2017 by Optimal Care Status (n=2412).
@@
We also explored the association between NSDoH profiles and the type of facility where patients were diagnosed or received treatment. Due to data limitations, we could not distinguish between these scenarios. For this sub-analysis, we dichotomized the facility type as academic medical center versus other types. Similarly to the optimal care analysis, we conducted a logistic regression adjusted for age, year, and insurance status at diagnosis, with results presented in Table B.2. Even after adjustment, NSDoH profiles 3,5, and 6 were significantly associated with receiving care at an academic medical center (vs. other). NSDoH profiles 3 and 6 were linked to higher odds of receiving care at academic centers, while profile 5 was associated with lower odds. The interactive map in 3.2 shows that profile 3 neighborhoods are often located near universities, which are in proximity to academic medical centers. In contrast, patients residing in neighborhoods in NSDoH profile 5 received care at either a community or teaching hospital (Table 2), and from observing the interactive map in 3.2 for NSDoH profile 5, most neighborhoods may be located farther from academic medical centers, making it less convenient for residents to access these facilities. Transportation challenges, despite lower probabilities of ”no vehicle in the household” in this profile, may still be a barrier, particularly for working-class populations (for which there is a high probability of 79%) with limited flexibility in work schedules.
Table B.2: Univariate and Multivariate Logistic Regression of Receiving Care at an Academic Medical Center Among Endometrial Cancer Cases Between 2015 and 2017 in the MCR (n=2412).
Model $1$Unadjusted model. Model $2$Adjusted for age, year, and insurance status at diagnosis.
@@
Profile 2 1.018 (0.792,1.311) 1.096 (0.856, 1.411)
Profile 3 1.439 (1.068,1.94) 1.452 (1.081, 1.954)
Profile 4 0.9 (0.683,1.178) 0.902 (0.679, 1.186)
Profile 5 0.577 (0.403,0.81) 0.588 (0.416, 0.823)
Profile 6 1.886 (1.325,2.671) 1.907 (1.299, 2.799)
Profile 7 0.959 (0.688,1.341) 1.001 (0.712, 1.413)
Profile 8 0.786 (0.545,1.131) 0.821 (0.561, 1.190)

### v2

@@ 2 Regression Analysis Additional
Figure 2.1: National Comprehensive Cancer Network guidelines example from 2020.
@@
Table 2.1 shows the distribution of patient characteristics by optimal care status defined as NCCN treatment adherence. For all the Bayesian logistic regression analyses, we used the brms package with default non-informative priors designed to have minimal influence on the results. The model was run with 2,000 iterations per chain, including a 500-iteration burn-in. Two chains were used, and they showed good mixing for all parameters, indicating no evidence of non-convergence.
Table 2.1: Sociodemographic and Clinical Characteristics of Massachusetts Cancer Registry Endometrial Cancer Cases from 2015-2017 by Optimal Care Status (n=2412).
@@
We also explored the association between NSDoH profiles and the type of facility where patients were diagnosed or received treatment. Due to data limitations, we could not distinguish between these scenarios. For this sub-analysis, we dichotomized the facility type as an academic medical center versus other types. Similarly to the optimal care analysis, we conducted a Bayesian logistic regression adjusted for age, year, and insurance status at diagnosis, with results in Table 2.2. Even after adjustment, NSDoH profiles 4 (racially/ethnically diverse (A+) and greater economic security and educational attainment) and 5 (racially/ethnically diverse (ABHL+), more renter-occupied housing with limited EN proficiency) were linked to higher odds of receiving care at academic centers. The interactive map in Section 3.2 shows that neighborhoods in profiles 4 and 5 are often located near universities, which are in proximity to academic medical centers. In contrast, although not statistically significant, patients residing in neighborhoods in NSDoH profile 3 (working class lower educational attainment) had lower odds of receiving care at an academic facility. These patients were more likely to receive care at either a community or teaching hospital (Table 2), and from observing the interactive map in Section 3.2 for NSDoH profile 3, most neighborhoods may be located farther from academic medical centers, making it less convenient for residents to access these facilities. Despite lower probabilities of ”no vehicle in the household” in this profile, transportation challenges may still be a barrier, particularly for working-class populations (for which there is a high probability of 79%) with limited flexibility in work schedules.
Table 2.2: Univariate and Multivariate Logistic Regression of Receiving Care at an Academic Medical Center Among Endometrial Cancer Cases Between 2015 and 2017 in the Massachusetts Cancer Registry (n=2412).
Model 1 Model 2
@@
Profile 2 1.049 (0.835, 1.324) 1.116 (0.875, 1.424)
Profile 3 0.860 (0.674, 1.083) 0.883 (0.702, 1.101)
Profile 4 1.701 (1.328, 2.181) 1.697 (1.313, 2.203)
Profile 5 2.011 (1.482, 2.730) 2.033 (1.464, 2.825)
  • Model 1: Unadjusted model.
  • Model 2: Adjusted for age, year, and insurance status at diagnosis.
  • Shorthand names for NSDoH Profiles: 1) advantaged non-Hispanic White, 2) disadvantaged racially/ethnically diverse (BHL+), more renter-occupied housing with limited EN proficiency, 3) working class, lower educational attainment, 4) racially/ethnically diverse (A+) and greater economic security and educational attainment, 5) racially/ethnically diverse (ABHL+), more renter-occupied housing with limited EN proficiency.